OpenStudy (anonymous):

PLEASE HELP! 3. Given f(x) = 5x + 1. What is f(3)? (Points : 4) 14 9 16 19 Question 4. 4. Which is a solution to the equation 6x + y = 4? (Points : 4) (1, 2) (1, 10) (2, 8) (2, 8) Question 5. 5. What is the slope of the graph of 8x 4y = 11? (Points : 4) Option A: 2 Option B: ½ Option C: ½ Option D: 2

OpenStudy (er.mohd.amir):

f(x)=5x+1 f(3)=5*3+1=15+1=16

OpenStudy (er.mohd.amir):

Q 4 ans (2,-8)

OpenStudy (er.mohd.amir):

if y=mx+c then m is slope write Eq in thhis form find slope